OpenStudy (anonymous):

How would you find the domain of the function: f(x)= (sqrt x+2) / (10-x) Must be in interval notation.

OpenStudy (tkhunny):

First, it's not clear. Is it \(\sqrt{x+2}\;or\;\sqrt{x}+2\) in the numerator?

OpenStudy (anonymous):

\[\sqrt{x+2}\]

OpenStudy (tkhunny):

Okay, start with "All Real Numbers" and throw out the bad guys. What Real Numbers don't work. What situations cause problems?

OpenStudy (anonymous):

\[\frac{ \sqrt{x+2} }{ 10-x }\]

OpenStudy (tkhunny):

Denominators can be a problem. Why? Square roots can be a problem. Why?